Bill leaves Histown at 11:30 am and drives towards Hertown, 86 miles away, at 55 mph. Sue leaves Hertown at 12:45 and drives towards Histown at 60 mph.
When and where will they meet?

Answer by stanbon(75887)

Bill leaves Histown at 11:30 am and drives towards Hertown, 86 miles away, at 55 mph. Sue leaves Hertown at 12:45 and drives towards Histown at 60 mph.
When and where will they meet?
-----
Bill DATA:
rate = 55 mph ; time = x hrs ; distance = rt = 55x miles
-------------------
Sue DATA:
rate = 60 mph ; time = (x-(5/4)) hrs ; distance = 60(x-(5/4)) miles
--------------------------------------------------------------
Equation:
distance + distance = 86 miles
55x + 60x - 5*15 = 86
115x = 161
x = 1.4 hrs (Time they meet will be 11:30+ 1.4 hrs = 12:54 pm
------
The meeting point will be 1.4(55) = 77 miles from Histown.
======================
Cheers,
Stan H.